OpenStudy (anonymous):

Armando read 20 pages of his 200-page book in 4 hours. At this rate, how long will it take him to read the entire book? 20 hours 40 hours 4 hours 5 hours

OpenStudy (anonymous):

What do you think it is first.

OpenStudy (anonymous):

20

OpenStudy (anonymous):

@Great_And_Powerful

OpenStudy (anonymous):

is it 20??

OpenStudy (anonymous):

why do you think its 20?

OpenStudy (anonymous):

because when u divide 200/4= 2

OpenStudy (anonymous):

ok, ok

OpenStudy (anonymous):

is it correct

OpenStudy (anonymous):

no, ok, 200/4=?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

no

OpenStudy (anonymous):

I would say 40.

OpenStudy (anonymous):

50

OpenStudy (anonymous):

leave hoslos plz

OpenStudy (anonymous):

Using ration and proportion, if 4h=20 pages, then xh=200pages, hence x= 200*4/20=40hours.

OpenStudy (anonymous):

yes hun, now 50\4

OpenStudy (anonymous):

Oh, my bad, my bad

OpenStudy (anonymous):

no, hold on

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

The message was arriving there, by the moment you said that. My bad!

OpenStudy (anonymous):

is ok

OpenStudy (anonymous):

go, you are giving direct answers

OpenStudy (anonymous):

200/20=?

OpenStudy (anonymous):

10

OpenStudy (anonymous):

ok, \[10\times4\]

OpenStudy (anonymous):

@Great_And_Powerful are u still here?? i an lagging over here

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

40

OpenStudy (anonymous):

yes!

OpenStudy (briellshaw):

you want the answer

OpenStudy (anonymous):

no

OpenStudy (anonymous):

it is 40 -_-

OpenStudy (anonymous):

ok thanks